OpenStudy (anonymous):

Solve. 3t < –15 (1 point) t < –5 t > –5 t < –45 t > –45

OpenStudy (anonymous):

divide each side by 3. t < -5

OpenStudy (anonymous):

can you help me with a couple more?

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

Solve. ≥ 12 (1 point) r ≤ 48 r ≥ 3 r ≥ 48 r ≤ 3

OpenStudy (anonymous):

@santw

OpenStudy (anonymous):

did you type that right? @SmithAshley

OpenStudy (anonymous):

Solve. r/4 ≥ 12 (1 point)r ≤ 48 r ≥ 3 r ≥ 48 r ≤ 3

OpenStudy (anonymous):

to get rid of the /4, multiply both sides by 4. after that you have \[4(\frac{r }{ 4 } \ge 12)\] the 4's cancel and leaves the 3rd option. \[r \ge48\]

OpenStudy (anonymous):

i'm trying to make sure you get the work too!

OpenStudy (anonymous):

can you help me with one more please?

OpenStudy (anonymous):

not the third, the second! ~_~ just post em, and notify me.

OpenStudy (anonymous):

Solve. –3y ≤ –18 (1 point) y ≤ –6 y ≤ 6 y ≥ 6 y ≤ 54 @santw

OpenStudy (anonymous):

divide each side by -3. so you get \[y \le6\]

OpenStudy (anonymous):

You only got 1 right and the other wrong :( @santw

OpenStudy (anonymous):

sorry! which one?

OpenStudy (anonymous):

number 2 and 3. Thats ok! :)

OpenStudy (anonymous):

3t < -15 t < -5 is right r/4>= 12 r >= 48 is also right -3y=< -18 however isn't y=< 6 because when you divide by a negative number you have to switch the sign around to make it y>= 6. Just for anyone finding this who may be confused by "SmithAshley" saying 2 were wrong when in fact only one was wrong.